# Intrinsic orbital fourfold anisotropic magnetoresistance in Dirac materials

### v1

@@ Abstract
To simulate the scatterings between surface Dirac electrons and the localized magnetic atoms (such as Mn atoms in $\mathrm{MnBi_{2}Te_{4}}$ (Wu et al., 2022)), we consider the spin-dependent scattering potential as $U(r)=\mu_{0}\sum_{i}\mathbf{M}\cdot\mathbf{\boldsymbol{\mu}}\delta(\mathbf{r}-\mathbf{R}_{i})$ (Chiba et al., 2017), where $\mu_{0}$ is the interaction strength, $\mathbf{\boldsymbol{\mu}}=\frac{1}{2}g\mu_{B}\mathbf{\boldsymbol{\sigma}}$ is spin magnetic moment of electrons, $R_{i}$ is the impurity position, and $\mathbf{M}$ is the field-polarized magnetic moment of the local impurities. We can calculate the relaxation time in the Born approximation and the electric conductivity of the surface states (SMA, ). The calculated conductivities for the surface and bulk states are shown in Fig. 2(a). The twofold AMR of the surface states can be understood from the distinction of the spin-momentum lockings. For the Dirac surface electrons, when the spins of electrons are parallel or antiparallel to the magnetic field or induced magnetization, the backscattering is forbidden, the corresponding resistivity will be much less affected by the magnetic field ($\rho_{\perp}$). On the other hand, when the spins are perpendicular to the field, the backscattering becomes allowed due to the broken time reversal symmetry, resulting in a significant enhancement of resistivity ($\rho_{\sslash}$). That is a positive anisotropic resistivity (i.e., $\rho_{\sslash}-\rho_{\perp}>0$). However, for the bulk states with parallel spin and momentum $\boldsymbol{k}\cdot\boldsymbol{\sigma}$, the direction of field enhanced resistivity and the direction of resistivity that is less sensitivitive to magnetic fields get exchanged, leading to a negative anisotropic resistivity ($\rho_{\sslash}-\rho_{\perp}<0$). It was also supported by the specific calculations of the anisotropic resistivity of a 2D slice of bulk states (SMA, )
@@
where $n_{s}$ is the density of magnetic impurities and $\theta=\mathrm{arctan}\;(M_{y}/M_{x})$. One can find that the cancellation of the twofold AMR between the surface states and the bulk states probably stems from the distinct spin-momentum lockings. In fact, this cancellation mechanism should be generally applicable to a wide range of 3D TIs including nonmagnetic, magnetic and higher-order ones (Xu et al., 2019; Zhang et al., 2019a, 2020).
@@
When the Fermi level lies in the conduction band of bulk, both the surface states and the bulk states simultaneously contribute to the electric transport. Because of the much hihger mobility, the surface states dominates over the bulk ones at lower fields. It is probable that, under some proper conditions, they can almost cancel out, making the fourfold AMR dominant. We numerically calculate the AMR that from both the bulk and the surface states of $\mathrm{MnBi_{2}Te_{4}}$ which is shown in Fig. 2(b). It can be seen that when the magnetic field is around 10 T, the fourfold AMR dominates. Further increasing of the magnetic field, the twofold AMR becomes dominant again, but the sign of its amplitude changes, which is in line with the observed AMR (Wu et al., 2022).
@@
Recently, the twofold AMR of Dirac surface states has been observed in the dual-gated devices of thin $\mathrm{Bi}{}_{2}\mathrm{SbTe}_{3}$ (Taskin et al., 2016; Zheng et al., 2020), where the Fermi level lies in the bulk gap. The electrical gating and chemical doping could tune the carrier density and shift the Fermi level into the bulk energy band, facilitating the realization of the predominant fourfold AMR therein.
@@
Summary.–We have found the tunable intrinsic orbital fourfold AMR due to OMM of electrons in various 3D Dirac materials. The unique spin-momentum lockings in topological bulk states and surface states cause competition between their twofold AMR, leaving predominant fourfold component. Our calculations are in quantitative agreement with the fourfold AMR in both bismuth and $\mathrm{MnBi_{2}Te_{4}}$ and the nontrivial magnetic-field dependence. This work reveals the significance of intrinsic OMM in novel AMR, inspiring more investigations of intriguing quantum phenomena associated with OMM in Dirac materials, such as the nonlinear PHE (He et al., 2019). Moreover, the first-principles calculations of OMM in the semiclassical transport theory could provide us a new and significant ingredient to deep investigate the novel magnetotransport properties of realistic quantum materials with complicated Fermi surfaces (Zhang et al., 2019b) and relaxation time (Dai et al., 2022).
@@
The authors acknowledge Y. Gao, Q. Niu, T. Qin and Z. Yuan for insightful discussions. This work was supported by the National Natural Science Foundation of China under Grants (No. U2032164 and No. 12174394) and by the High Magnetic Field Laboratory of Anhui Province and by the HFIPS Director’s Fund (Grant Nos. YZJJQY202304 and BJPY2023B05) and Chinese Academy of Sciences under contract No. JZHKYPT-2021-08.

### v2

@@ Abstract
To simulate the scatterings between surface Dirac electrons and the localized magnetic atoms (such as Mn atoms in $\mathrm{MnBi_{2}Te_{4}}$ [33]), we consider the spin-dependent scattering potential as $U(r)=\mu_{0}\sum_{i}\mathbf{M}\cdot\mathbf{\bm{\mu}}\delta(\mathbf{r}-\mathbf{R}_{i})$, similar to the case of topological insulator-ferromagnetic insulator bilayer [75], where $\mu_{0}$ is the interaction strength, $\mathbf{\bm{\mu}}=\frac{1}{2}g\mu_{B}\mathbf{\bm{\sigma}}$ is spin magnetic moment of electrons, $R_{i}$ is the impurity position, and $\mathbf{M}$ is the field-polarized magnetic moment of the local impurities. We can calculate the relaxation time in the Born approximation and the electric conductivity of the surface states [50]. The calculated conductivities for the surface and bulk states are shown in Fig. 2(a). The twofold AMR of the surface states can be understood from the distinction of the spin-momentum lockings. For the Dirac surface electrons, when the spins of electrons are parallel or antiparallel to the magnetic field or induced magnetization, the backscattering is forbidden, the corresponding resistivity will be much less affected by the magnetic field ($\rho_{\perp}$). On the other hand, when the spins are perpendicular to the field, the backscattering becomes allowed due to the broken time reversal symmetry, resulting in a significant enhancement of resistivity ($\rho_{\parallel}$). That is a positive anisotropic resistivity (i.e., $\rho_{\parallel}-\rho_{\perp}>0$). However, for the bulk states with parallel spin and momentum $\bm{k}\cdot\bm{\sigma}$, the direction of field enhanced resistivity and the direction of resistivity that is less sensitive to magnetic fields get exchanged, leading to a negative anisotropic resistivity ($\rho_{\parallel}-\rho_{\perp}<0$). It was also supported by the specific calculations of the anisotropic resistivity of a 2D slice of bulk states [50]
@@
where $n_{s}$ is the density of magnetic impurities and $\theta=\mathrm{arctan}\;(M_{y}/M_{x})$. One can find that the cancellation of the twofold AMR between the surface states and the bulk states probably stems from the distinct spin-momentum lockings, highlighting the role of the interplay between the magnetic scatterings and SOC in AMR [76]. In fact, this cancellation mechanism should be generally applicable to a wide range of 3D TIs including nonmagnetic, magnetic and higher-order ones [77, 78, 79].
@@
When the Fermi level lies in the conduction band of bulk, both the surface states and the bulk states simultaneously contribute to the electric transport. Because of the much higher mobility, the surface states dominates over the bulk ones at lower fields. It is probable that, under some proper conditions, they can almost cancel out, making the fourfold AMR dominant. We numerically calculate the AMR that from both the bulk and the surface states of $\mathrm{MnBi_{2}Te_{4}}$, as shown in Fig. 2(b). It can be seen that when the magnetic field is around 10 T, the orbital fourfold AMR dominates. Further increasing of the magnetic field, the orbital twofold AMR becomes dominant again, but the sign of its amplitude changes, which is in line with the observed AMR [33].
@@
Recently, when the Fermi level lies in the bulk gap, the twofold AMR of Dirac surface states has been observed in the dual-gated devices of thin films of nonmagnetic TI $\mathrm{Bi}{}_{2}\mathrm{SbTe}_{3}$ [80] and been understood in different mechanisms, such as scatterings of magnetic impurities [80] and tilting of surface Dirac cones with nonlinear momentum terms [81]. The electrical gating and chemical doping could tune the carrier density and shift the Fermi level into the bulk energy band, facilitating the realization of the predominant orbital fourfold AMR therein.
@@
Summary.–We have found the tunable orbital fourfold AMR due to OMM of electrons in various 3D Dirac materials. The unique spin-momentum lockings in topological bulk states and surface states cause competition between their twofold AMR, leaving predominant fourfold component. Our calculations are in quantitative agreement with the fourfold AMR in both bismuth and $\mathrm{MnBi_{2}Te_{4}}$ and the nontrivial magnetic-field dependence. This work reveals the significance of intrinsic OMM in novel AMR, inspiring more investigations of intriguing quantum phenomena associated with OMM in quantum materials, such as the nonlinear PHE [82]. Moreover, the first-principles calculations of OMM in the semiclassical transport theory could provide us a new and significant ingredient to deep investigate the novel magnetotransport properties of realistic quantum materials with complicated Fermi surfaces [83] and relaxation time [20].
@@
The authors acknowledge Y. Gao, Q. Niu, T. Qin and Z. Yuan for insightful discussions. This work was supported by National Key R&D Program of the MOST of China (Grant No. 2024YFA1611300) and the National Natural Science Foundation of China under Grants (No. U2032164 and No. 12174394) and by the High Magnetic Field Laboratory of Anhui Province and by the HFIPS Director’s Fund (Grant Nos. YZJJQY202304 and BJPY2023B05) and Chinese Academy of Sciences under contract No. JZHKYPT-2021-08.